OpenStudy (anonymous):

The equation of the line parallel to the line with equation y= -1/2x and containing the point (1,3) is_______. a. y=1/2x + 5/2 b. y= -2x+5 c. y=-1/2x+7/2 d. y=-1/2x+2

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Well, it needs to have the same slope, -1/2, so that rules out a and b. Then it needs to pass through point (1, 3), so plug 1 into both c and d and see if you get 3: c) y = -1/2 * 1 + 7/2 = 3 (so c it is) d) y = -1/2 * 1 + 2 = 1.5 (nope)